# Open-Set Domain Adaptation Under Background Distribution Shift: Challenges and A Provably Efficient Solution

### v2

@@ A.1 Theoretical analysis
${\mathbf{w}}_{\mathrm{color}}^{\top}(y_{i}\mu+y_{i}\xi_{i})$ $={\mathbf{w}}_{\mathrm{color}}^{\top}(\mu+\xi_{i})$
$=\left(\frac{\alpha r^{2}_{\eta}}{r^{2}_{\mu}}-\frac{5\tau}{r^{2}_{\mu}}\right)(\mu^{\top}\mu+\mu^{\top}\xi_{i})-\beta(\mu^{\top}\bar{\xi}_{{\mathcal{T}},0}+\xi_{i}^{\top}\bar{\xi}_{{\mathcal{T}},0})$
We know that $|\mu^{\top}\xi_{i}|\leq\sigma tr_{\mu}$ w.p. at least $1-2\exp\left(-ct^{2}/2\right)$ and $|\xi_{i}^{\top}\bar{\xi}_{{\mathcal{T}},0}|\leq\sigma^{2}t\sqrt{\frac{d}{|U|-k}}$ w.p. at least $1-2\exp(-ct^{2})$ for $t\leq\sigma^{2}$ and $t\leq\frac{\sqrt{d}}{2(2+\sqrt{|U|-1})}$. We also know that $|\mu^{\top}\bar{\xi}_{{\mathcal{T}},0}|\leq\frac{\sigma tr_{\mu}}{|U|-k}$ w.p. at least $1-2\exp(-ct^{2}/2)$ Hence, w.p. at least $1-4\exp\left(-ct^{2}/2\right)-2\exp(-ct^{2})$, we can say that,
${\mathbf{w}}_{\mathrm{color}}^{\top}(-y_{i}\mu+y_{i}\xi_{i})$ $\geq\left(\frac{\alpha r^{2}_{\eta}}{r^{2}_{\mu}}-\frac{5\tau}{r^{2}_{\mu}}\right)(r^{2}_{\mu}-\sigma tr_{\mu})-\beta\left(\frac{\sigma tr_{\mu}}{|U|-k}+\sigma^{2}t\sqrt{\frac{d}{|U|-k}}\right)$
$\geq\frac{r^{2}_{\mu}}{2}\left(\frac{\alpha r^{2}_{\eta}}{r^{2}_{\mu}}-\frac{5\tau}{r^{2}_{\mu}}\right)-\beta\left(\frac{\sigma tr_{\mu}}{|U|-k}+\sigma^{2}t\sqrt{\frac{d}{|U|-k}}\right)\text{ ..... for $r_{\mu}\geq 2\sigma t$}$
@@
We want ${\mathbf{w}}_{\mathrm{color}}^{\top}(\mu-\eta)\leq-\tau$ i.e. ${\mathbf{w}}_{\mathrm{color}}^{\top}(\eta-\mu)\geq\tau$
@@
${\mathbf{w}}_{\mathrm{color}}^{\top}(\eta-\mu)$ $=\alpha\eta^{\top}\eta-\left(\frac{\alpha r^{2}_{\eta}}{r^{2}_{\mu}}-\frac{5\tau}{r^{2}_{\mu}}\right)\mu^{\top}\mu+\beta\cdot\mu^{\top}\bar{\xi}_{{\mathcal{T}},0}$
We know that $|\mu^{\top}\bar{\xi}_{{\mathcal{T}},0}|\leq\frac{\sigma tr_{\mu}}{|U|-k}$ w.p. at least $1-2\exp(-ct^{2}/2)$ Hence, w.p. at least $1-2\exp\left(-ct^{2}/2\right)$, we can say that,
${\mathbf{w}}_{\mathrm{color}}^{\top}(\eta-\mu)$ $\geq\alpha r^{2}_{\eta}-\left(\frac{\alpha r^{2}_{\eta}}{r^{2}_{\mu}}-\frac{5\tau}{r^{2}_{\mu}}\right)r^{2}_{\mu}-\beta\frac{\sigma tr_{\mu}}{|U|-k}$
@@
${\mathbf{w}}_{\mathrm{color}}=3\frac{\eta}{r^{2}_{\eta}}+\left(\frac{3}{r^{2}_{\mu}}-\frac{5\tau}{r^{2}_{\mu}}\right)\mu-\sigma^{-2}\frac{\sqrt{|U|-k}}{d}\bar{\xi}_{S_{{\mathcal{T}},0}}$
@@
$\leq\frac{3}{8}\frac{1-\tau}{\sqrt{(|U|)^{-1}\sigma^{2}d}}....\text{for $t\leq\frac{3}{2}\sqrt{d}$}$
$\leq\frac{3}{8}(1-\tau)\sqrt{|U|}....\text{for $\sigma=\frac{1}{\sqrt{d}}$}$
@@
Consider a Linear-Gaussian PU-learning problem with parameters $\mu,\eta,d>300,\alpha\in{(0,\frac{1}{1024})}$ and dataset sizes $_{{\mathcal{T}}}>10,N_{{\mathcal{S}}}=N_{{\mathcal{T}}}, N=N$$N_{{\mathcal{S}}}+N_{{\mathcal{T}}}=2N_{{\mathcal{T}}}$, and let $\delta\in{(0,1)}$. For all problems where
$\text{(i)}\;\min\{r_{\eta},r_{\mu}\}\geq\tfrac{32}{\sqrt{N}},\quad\text{(ii)}\;r_{\mu}\leq\tfrac{1}{2}\sqrt{N_{{\mathcal{T}},0}},\quad\text{(iii)}\;\tfrac{r_{\eta}}{r_{\mu}}\leq\tfrac{32}{N},$ (8)
$\text{(iv)}\;d>c_{1}N\log\!\left(\tfrac{c_{2}}{\delta}\right),\quad\text{(v)}\;N\geq c_{1}\log\!\left(\tfrac{c_{2}}{\delta}\right).$ (9)
@@
Based on the stated assumptions on $N_{\mathcal{S}},N_{{\mathcal{T}}},N_{{\mathcal{T}},0},N_{{\mathcal{T}},1}$
v) $\sqrt{N\cdot N_{{\mathcal{T}},1}}=\sqrt{2N_{{\mathcal{T}}}\cdot N_{{\mathcal{T}},1}}\geq\sqrt{2N_{{\mathcal{T}},0}\cdot N_{{\mathcal{T}},1}}\geq(\sqrt{2}-1)\sqrt{N_{{\mathcal{T}},0}\cdot N_{{\mathcal{T}},1}}+\sqrt{N_{{\mathcal{T}},0}\cdot N_{{\mathcal{T}},1}}\geq 1+\sqrt{N_{{\mathcal{T}},0}\cdot N_{{\mathcal{T}},1}}$
Using condition ii) we can write, $1+\sqrt{N_{{\mathcal{T}},0}\cdot N_{{\mathcal{T}},1}}\geq 1+2r_{\mu}\sqrt{N_{{\mathcal{T}},1}}$
vi) $r_{\mu}\sqrt{N}\geq 8+\frac{1}{2}r_{\mu}\sqrt{N}\geq 8+16r_{\mu}\sqrt{N_{{\mathcal{T}},1}}=8(1+2r_{\mu}\sqrt{N_{{\mathcal{T}},1}})$. Here, the first transition is from condition i) and second from the range of $\alpha=\frac{N_{{\mathcal{T}},1}}{N_{{\mathcal{T}}}}\leq\frac{1}{1024}$.
@@
$N$ $\geq c_{1}\log(\frac{c_{2}}{\delta})$
$d$ $\geq c_{1}N\log(\frac{c_{2}}{\delta})$
$\frac{1}{2\sqrt{N}}$ $\leq\min(r_{\eta},r_{\mu})$
$r_{\mu}\$$\leq\sqrt{N\cdot N_{{\mathcal{T}},0}}$
$\frac{r_{\eta}}{r_{\mu}}$ $\leq\frac{1}{4}\left(\frac{1}{1+2r_{\mu}\sqrt{N_{{\mathcal{T}},1}}}-\frac{4}{r_{\mu}\sqrt{N}}\right)$
$\frac{3}{r_{\eta}}+\frac{3-5\tau}{r_{\mu}}$ $\leq\frac{27}{80}\sqrt{N}-2$
Then putting this together, for parameters that satisfy the above constraint we have from i) and iv),
$\frac{r_{\eta}}{r_{\mu}}$ $<\frac{1}{8}(N\cdot N_{{\mathcal{T}},1})^{-1/2}$
$<\frac{1}{8}(1+2r_{\mu}\sqrt{N_{{\mathcal{T}},1}})^{-1}\text{ \hskip 5.69054pt ... from v)}$
$<\frac{1}{4}\left(\frac{1}{1+2r_{\mu}\sqrt{N_{{\mathcal{T}},1}}}-\frac{4}{r_{\mu}\sqrt{N}}\right)\text{ \hskip 5.69054pt ... from vi)}$
$\frac{1}{2\sqrt{N}}$ $\leq r_{\mu}<\min\left(\frac{1}{2}\sqrt{d}(1-5\tau)\sqrt{N_{{\mathcal{T}},0}},4\sqrt{d}\tau\sqrt{N_{{\mathcal{T}},0}}\right),$
$<\frac{1}{4}\sqrt{dN_{{\mathcal{T}},0}}\textbf{\hskip 5.69054pt ... using $\tau=0.1$}$
$\frac{3}{r_{\eta}}+\frac{3-5\tau}{r_{\mu}}$ $\leq\frac{3}{8}(1-\tau)\sqrt{N}-2$
$\frac{3}{r_{\eta}}+\frac{2.5}{r_{\mu}}$ $\leq\frac{27}{80}\sqrt{N}-2\textbf{\hskip 5.69054pt ... using $\tau=0.1$}$
$3+\frac{2.5r_{\eta}}{r_{\mu}}$ $\leq\frac{27}{80}r_{\eta}(\sqrt{N}-2)$
$\frac{r_{\eta}}{r_{\mu}}\leq\frac{10}{25}\left(\frac{27}{80}r_{\eta}(\sqrt{N}-2)-3\right)$
We know from ii) that $r_{\eta}(\sqrt{N}-2)\geq 16\implies\frac{27}{5}\leq\frac{27}{80}r_{\eta}(\sqrt{N}-2)$.
@@
$\leq\frac{10}{25}\left(\frac{27}{80}r_{\eta}(\sqrt{N}-2)-3\right)$
@@ A.2 Concentration statements
  • Bound on norm of vector (ledoux2013probability, Eq. 3.5)
@@
    This is a derivation from Bernstein’s inequality, see e.g. Lemma 3 in puli2023don.
@@
    Recall Bernstein (vershynin2018high) says that for independent sub-exponential variables $X_{1},\ldots,X_{N}$, for every $t>0$,
@@
    Taking union bound of inequalities 13 and 14, we get
@@ A.3 Proof of Lemma 1
Let us recall the strong positivity assumption stated in the main paper, which appears in garg22adaptation.
@@ A.4.1 AUROC vs AUPRC scores for novel class detection
It is a common argument in the machine learning literature that AUPRC scores are more suitable for evaluating methodologies in class-imbalanced scenarios. However, this stance is nuanced, as some research, such as mcdermott2024closer suggests favoring AUROC over AUPRC in certain imbalanced conditions. mcdermott2024closer further notes that AUPRC inherently emphasizes the performance on samples with higher scores. Given our experimental focus on assessing models’ ability to assign higher scores for novelty detection, AUPRC emerges as the most relevant metric for our analysis, especially when the proportions of positive class (novel class in our case) is very low.
@@ A.5.1 Dataset
CIFAR100: It consists of 60,000 32x32 colour images in 20 primary classes (superclasses) each of which have 5 subcategories composing a total of 100 classes krizhevsky2009learning. The training set has 50,000 images (i.e. 500 images per subcategory and 2500 images per primary class) while the test set has 10000 images (i.e. 100 images per subcategory and 500 images per primary class). We retain these splits for our experiments. We use 4 primary categories as aquatic mammals, flowers, fishes and birds. The subcategories of the aquatic mammals are beaver, dolphin, otter, seal are whale while that of flowers are orchid, poppy, rose, sunflower and tulip. Similarly fishes and birds have 5 subcategories each. We vary the marginal distribution of these subcategories to create a subpopulation shift leading to a background shift between source and target data while maintaining no label shift w.r.t. primary categories i.e. aquatic mammals and flowers. The novel category is randomly selected from the remaining unseen categories.

### v3

@@ A.1 Theoretical analysis
${\mathbf{w}}_{\mathrm{color}}^{\top}(y_{i}\bm{\mu}+y_{i}\xi_{i})$ $={\mathbf{w}}_{\mathrm{color}}^{\top}(\bm{\mu}+\xi_{i})$
$=\left(\frac{\alpha r^{2}_{\eta}}{r^{2}_{\mu}}-\frac{5\tau}{r^{2}_{\mu}}\right)(\bm{\mu}^{\top}\bm{\mu}+\bm{\mu}^{\top}\xi_{i})-\beta(\bm{\mu}^{\top}\bar{\xi}_{{\mathcal{T}},0}+\xi_{i}^{\top}\bar{\xi}_{{\mathcal{T}},0})$
We know that $|\bm{\mu}^{\top}\xi_{i}|\leq\sigma tr_{\mu}$ w.p. at least $1-2\exp\left(-ct^{2}/2\right)$ and $|\xi_{i}^{\top}\bar{\xi}_{{\mathcal{T}},0}|\leq\sigma^{2}t\sqrt{\frac{d}{|U|-k}}$ w.p. at least $1-2\exp(-ct^{2})$ for $t\leq\sigma^{2}$ and $t\leq\frac{\sqrt{d}}{2(2+\sqrt{|U|-1})}$. We also know that $|\bm{\mu}^{\top}\bar{\xi}_{{\mathcal{T}},0}|\leq\frac{\sigma tr_{\mu}}{|U|-k}$ w.p. at least $1-2\exp(-ct^{2}/2)$ Hence, w.p. at least $1-4\exp\left(-ct^{2}/2\right)-2\exp(-ct^{2})$, we can say that,
${\mathbf{w}}_{\mathrm{color}}^{\top}(-y_{i}\bm{\mu}+y_{i}\xi_{i})$ $\geq\left(\frac{\alpha r^{2}_{\eta}}{r^{2}_{\mu}}-\frac{5\tau}{r^{2}_{\mu}}\right)(r^{2}_{\mu}-\sigma tr_{\mu})-\beta\left(\frac{\sigma tr_{\mu}}{|U|-k}+\sigma^{2}t\sqrt{\frac{d}{|U|-k}}\right)$
$\geq\frac{r^{2}_{\mu}}{2}\left(\frac{\alpha r^{2}_{\eta}}{r^{2}_{\mu}}-\frac{5\tau}{r^{2}_{\mu}}\right)-\beta\left(\frac{\sigma tr_{\mu}}{|U|-k}+\sigma^{2}t\sqrt{\frac{d}{|U|-k}}\right)\text{ \hskip 11.38109pt for $r_{\mu}\geq 2\sigma t$}$
@@
We want ${\mathbf{w}}_{\mathrm{color}}^{\top}(\bm{\mu}-\bm{\eta})\leq-\tau$ i.e. ${\mathbf{w}}_{\mathrm{color}}^{\top}(\bm{\eta}-\bm{\mu})\geq\tau$
@@
${\mathbf{w}}_{\mathrm{color}}^{\top}(\bm{\eta}-\bm{\mu})$ $=\alpha\bm{\eta}^{\top}\bm{\eta}-\left(\frac{\alpha r^{2}_{\eta}}{r^{2}_{\mu}}-\frac{5\tau}{r^{2}_{\mu}}\right)\bm{\mu}^{\top}\bm{\mu}+\beta\cdot\bm{\mu}^{\top}\bar{\xi}_{{\mathcal{T}},0}$
We know that $|\bm{\mu}^{\top}\bar{\xi}_{{\mathcal{T}},0}|\leq\frac{\sigma tr_{\mu}}{|U|-k}$ w.p. at least $1-2\exp(-ct^{2}/2)$ Hence, w.p. at least $1-2\exp\left(-ct^{2}/2\right)$, we can say that,
${\mathbf{w}}_{\mathrm{color}}^{\top}(\bm{\eta}-\bm{\mu})$ $\geq\alpha r^{2}_{\eta}-\left(\frac{\alpha r^{2}_{\eta}}{r^{2}_{\mu}}-\frac{5\tau}{r^{2}_{\mu}}\right)r^{2}_{\mu}-\beta\frac{\sigma tr_{\mu}}{|U|-k}$
@@
${\mathbf{w}}_{\mathrm{color}}=3\frac{\bm{\eta}}{r^{2}_{\eta}}+\left(\frac{3}{r^{2}_{\mu}}-\frac{5\tau}{r^{2}_{\mu}}\right)\bm{\mu}-\sigma^{-2}\frac{\sqrt{|U|-k}}{d}\bar{\xi}_{S_{{\mathcal{T}},0}}$
@@
$\leq\frac{3}{8}\frac{1-\tau}{\sqrt{(|U|)^{-1}\sigma^{2}d}}\text{ \hskip 11.38109pt for $t\leq\frac{3}{2}\sqrt{d}$}$
$\leq\frac{3}{8}(1-\tau)\sqrt{|U|}\text{\hskip 11.38109pt for $\sigma=\frac{1}{\sqrt{d}}$}$
@@
Consider a Linear-Gaussian PU-learning problem with parameters $\bm{\mu},\bm{\eta},d>300$, dataset sizes $_{{\mathcal{T}}}>10, N=N$$N_{{\mathcal{S}}}+N_{{\mathcal{T}}}$, $\alpha\in{(0,\frac{N}{1024N_{\mathcal{T}}})}$ and let $\delta\in{(0,1)}$. For all problems where
$\min\{r_{\eta},r_{\mu}\}$ $\geq\tfrac{16}{\sqrt{N_{\mathcal{T}}}},$ (8)
$r_{\mu}$ $\leq\tfrac{1}{2}\sqrt{N_{{\mathcal{T}},0}},$ (9)
$\tfrac{r_{\eta}}{r_{\mu}}$ $\leq\tfrac{4}{N},$ (10)
$c_{1}\log\!\left(\tfrac{c_{2}}{\delta}\right)$ $\leq\min{\left(\sqrt{N},\sqrt{\frac{d}{N}}\right)}$ (11)
@@
Let us denote $p=\frac{N}{N_{\mathcal{T}}}$. Then based on the stated assumptions on $N_{\mathcal{S}},N_{{\mathcal{T}}},N_{{\mathcal{T}},0},N_{{\mathcal{T}},1}$ we can say,
$\sqrt{N\cdot N_{{\mathcal{T}},1}}$ $=\sqrt{pN_{{\mathcal{T}}}\cdot N_{{\mathcal{T}},1}}$
$\geq\sqrt{pN_{{\mathcal{T}},0}\cdot N_{{\mathcal{T}},1}}$
$\geq(\sqrt{p}-1)\sqrt{N_{{\mathcal{T}},0}\cdot N_{{\mathcal{T}},1}}+\sqrt{N_{{\mathcal{T}},0}\cdot N_{{\mathcal{T}},1}}$
$\geq 1+\sqrt{N_{{\mathcal{T}},0}\cdot N_{{\mathcal{T}},1}}$ (12)
Using eq.˜9 we can write, $1+\sqrt{N_{{\mathcal{T}},0}\cdot N_{{\mathcal{T}},1}}\geq 1+2r_{\mu}\sqrt{N_{{\mathcal{T}},1}}$
From eq.˜8 we know that $r_{\mu}\sqrt{N}\geq 16$. Hence,
$r_{\mu}\sqrt{N}$ $\geq 8+\frac{1}{2}r_{\mu}\sqrt{N}$
$\geq 8+\frac{1}{2}r_{\mu}\sqrt{\frac{p}{\alpha}\cdot N_{{\mathcal{T}},1}}$
$\geq 8+16r_{\mu}\sqrt{N_{{\mathcal{T}},1}}=8(1+2r_{\mu}\sqrt{N_{{\mathcal{T}},1}})$ (13)
Here, the first transition is from eq.˜8 while second and third from the range of $\alpha=\frac{pN_{{\mathcal{T}},1}}{N}\leq\frac{p}{1024}$.
@@
$\sqrt{N}$ $\geq c_{1}\log(\frac{c_{2}}{\delta})$ (14)
$\sqrt{d}$ $\geq c_{1}\sqrt{N}\log(\frac{c_{2}}{\delta})$ (15)
[…]
$\frac{r_{\eta}}{r_{\mu}}$ $\leq\frac{1}{4}\left(\frac{1}{1+2r_{\mu}\sqrt{N_{{\mathcal{T}},1}}}-\frac{4}{r_{\mu}\sqrt{N}}\right)$ (18)
$\frac{3}{r_{\eta}}+\frac{3-5\tau}{r_{\mu}}$ $\leq\frac{27}{80}\sqrt{N}-2$ (19)
Then putting this together, for parameters that satisfy the eq.˜11 we have,
$\frac{r_{\eta}}{r_{\mu}}$ $\leq\frac{4}{N}=4\cdot\sqrt{\frac{\alpha}{N\cdot N_{{\mathcal{T}},1}p}}$
$\leq\frac{1}{8}(N\cdot N_{{\mathcal{T}},1})^{-1/2}\text{ \hskip 11.38109pt as $\alpha\leq\frac{p}{1024}$}$
[…]
which satisfies the constraint in proposition 1.
$\frac{1}{2\sqrt{N_{\mathcal{T}}}}$ $\leq r_{\mu}<\min\left(\frac{1}{2}\sqrt{d}(1-5\tau)\sqrt{N_{{\mathcal{T}},0}},4\sqrt{d}\tau\sqrt{N_{{\mathcal{T}},0}}\right),$
$<\frac{1}{4}\sqrt{dN_{{\mathcal{T}},0}}\text{\hskip 11.38109pt using $\tau=0.1$}$
$\frac{3}{r_{\eta}}+\frac{3-5\tau}{r_{\mu}}$ $\leq\frac{3}{8}(1-\tau)\sqrt{N_{\mathcal{T}}}-2$
$\frac{3}{r_{\eta}}+\frac{2.5}{r_{\mu}}$ $\leq\frac{27}{80}\sqrt{N_{\mathcal{T}}}-2\text{\hskip 11.38109pt using $\tau=0.1$}$
$3+\frac{2.5r_{\eta}}{r_{\mu}}$ $\leq\frac{27}{80}r_{\eta}(\sqrt{N_{\mathcal{T}}}-2)$
$\frac{r_{\eta}}{r_{\mu}}\leq\frac{10}{25}\left(\frac{27}{80}r_{\eta}(\sqrt{N_{\mathcal{T}}}-2)-3\right)$
We know from eq.˜8 that $r_{\eta}(\sqrt{N_{\mathcal{T}}}-2)\geq 16\implies\frac{27}{5}\leq\frac{27}{80}r_{\eta}(\sqrt{N_{\mathcal{T}}}-2)$.
@@
$\leq\frac{10}{25}\left(\frac{27}{80}r_{\eta}(\sqrt{N_{\mathcal{T}}}-2)-3\right)$
This means $\frac{r_{\eta}}{r_{\mu}}\leq\min{(\frac{24}{25},\frac{4}{N_{\mathcal{T}}})}$ which satisfies eq.˜10. ∎
@@ A.2 Concentration statements
  • Bound on norm of vector (Ledoux and Talagrand, 2013, Eq. 3.5)
@@
    This is a derivation from Bernstein’s inequality, see e.g. Lemma 3 in Puli et al. (2023).
@@
    Recall Bernstein (Vershynin, 2018) says that for independent sub-exponential variables $X_{1},\ldots,X_{N}$, for every $t>0$,
@@
    Taking union bound of inequalities 23 and 24, we get
@@ A.3 Proof of Lemma 1
Let us recall the strong positivity assumption stated in the main paper, which appears in Garg et al. (2022).
@@ A.4.1 AUROC vs AUPRC scores for novel class detection
It is a common argument in the machine learning literature that AUPRC scores are more suitable for evaluating methodologies in class-imbalanced scenarios. However, this stance is nuanced, as some research, such as McDermott et al. (2024) suggests favoring AUROC over AUPRC in certain imbalanced conditions. McDermott et al. (2024) further notes that AUPRC inherently emphasizes the performance on samples with higher scores. Given our experimental focus on assessing models’ ability to assign higher scores for novelty detection, AUPRC emerges as the most relevant metric for our analysis, especially when the proportions of positive class (novel class in our case) is very low.
@@ A.5.1 Dataset
CIFAR100: It consists of 60,000 32x32 colour images in 20 primary classes (superclasses) each of which have 5 subcategories composing a total of 100 classes Krizhevsky et al. (2009). The training set has 50,000 images (i.e. 500 images per subcategory and 2500 images per primary class) while the test set has 10000 images (i.e. 100 images per subcategory and 500 images per primary class). We retain these splits for our experiments. We use 4 primary categories as aquatic mammals, flowers, fishes and birds. The subcategories of the aquatic mammals are beaver, dolphin, otter, seal are whale while that of flowers are orchid, poppy, rose, sunflower and tulip. Similarly fishes and birds have 5 subcategories each. We vary the marginal distribution of these subcategories to create a subpopulation shift leading to a background shift between source and target data while maintaining no label shift w.r.t. primary categories i.e. aquatic mammals and flowers. The novel category is randomly selected from the remaining unseen categories.